OpenStudy (anonymous):

Hello could someone please tell me which trigonometric identities do I need to apply in order to get this result?

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi}\sqrt{1-\cos \Theta}d\] the answer to simplification : \[\int\limits_{0}^{2\pi}\sqrt{2\sin^2 \frac{ \theta}{ 2 }}d\]